mialopez:

The equation y minus 7 Over 2 = one-half (x minus 4) is written in point-slope form. What is the y-intercept of the equation?

rf6086:

none

mialopez:

nope

mialopez:

Would you like me to give you the answer choices?

rf6086:

yeah

mialopez:

OK

mialopez:

Negative 11 Over 2 Negative one-half One-half Three-halves

mialopez:

A: -11/2 B: - 1/2 C: 1/2 D: 3/2

loveboy:

y-7/2 = 1/2(x-4) Plug x = 0, y-7/2 = -2 y = 3/2 Option D would be correct